Given the demand function for the two markets and total cost function that:

Q= 55 – P1; Q= 70 – 2P2 and TC=5Q + 20 where Q = Q1 + Q2

A.   Calculate the equilibrium price, equilibrium quantity in each market with price discrimination.

B.    Calculate the maximum profit using equilibrium price and quantity with price discrimination.

C.    Calculate the price elasticity of demand in each market using MR=P(1+1/Ɛp).

D.   Calculate the equilibrium price and quantity without price discrimination.

E.    Calculate the maximum profit without price discrimination.

F.     Is the Monopolist better off with or without price discrimination.



1
Expert's answer
2022-01-12T08:34:55-0500

Solution:

A.). Derive the inverse demands for both markets:

Market 1:

Q1 = 55 – P1

P1 = 55 – Q1

TR1 = P1 "\times" Q1 = (55 – Q1) x Q1 = 55Q1 – Q12

Derive MR1 from TR1:

MR1 = "\frac{\partial TR_{1} } {\partial Q_{1} }" = 55 – 2Q1

Market 2:

Q2 = 70 – 2P2

P2 = 35 – 0.5Q2

TR2 = P2 "\times" Q2 = (70 – Q2) "\times" Q2 = 70Q2 – Q22

Derive MR2 from TR2:

MR2 = "\frac{\partial TR_{2} } {\partial Q_{2} }" = 70 – 2Q2

 

Derive MC:

MC = "\frac{\partial TC } {\partial Q }" = 5

 

Set MR1 = MC

55 – 2Q1 = 5

50 = 2Q1

Q1 = 25

Substitute to get P1:

P1 = 55 – Q1 = 55 – 25 = 30

Market 1 equilibrium quantity (Q1) = 30

Market 1 equilibrium Price (P1) = 25

 

Set MR2 = MC

70 – 2Q2 = 5

65 = 2Q2

Q2 = 32.5

Substitute to get P2:

P2 = 35 – 0.5Q2 = 35 – 0.5(32.5) = 35 – 16.25 = 18.75

Market 2 equilibrium quantity (Q1) = 32.5

Market 2 equilibrium Price (P1) = 18.75

 

B.). Profit Market 1 = TR1 – TC

TR1 = P1 "\times" Q1 = 30 "\times" 25 = 750

TC = 5Q + 20 = 5(30 + 32.5) + 20 = 312.5 + 20 = 332.5

Profit Market 1 = 750 – 332.5 = 437.5

 

Profit Market 2 = TR2 – TC

TR2 = P2 "\times" Q2 = 32.5 "\times" 18.75 = 609.38

TC = 5Q + 20 = 5(30 + 32.5) + 20 = 312.5 + 20 = 332.5

Profit Market 2 = 609.38 – 332.5 = 276.88

 

C.). Price elasticity Market 1 = "\frac{\triangle Q_{1} }{\triangle P_{1}} \times \frac{ Q_{1} }{P_{1}}"

"\frac{\triangle Q_{1} }{\triangle P_{1}}" = -1

= -1 "\times \frac{25}{30} = -0.83"

Price elasticity Market 1 = -0.83

 

Price elasticity Market 2 = "\frac{\triangle Q_{2} }{\triangle P_{2}} \times \frac{ Q_{2} }{P_{2}}"

"\frac{\triangle Q_{2} }{\triangle P_{2}}" = -2

= -2 "\times" "\frac{18.75}{32.5} = -1.15"

Price elasticity Market 2 = -1.15